# In-Context Reinforcement Learning for Variable Action Spaces

### v3

@@ 5 Ablations
Action Set Prompt: We assess the impact of omitting the action embedding enumeration from the model context. The tuned hyperparameters can be found in Table 4.
@@
Figure 7(a) reveals that removing the action set prompt does not significantly alter the number of unique actions attempted by each model. This suggests that models without the prompt continue to sample a diverse range of actions, comparable to the original setup. However, as shown in Figure 7(b), the presence of the prompt does affect performance. We hypothesize that, without the action set prompt, the model lacks explicit knowledge of the action space, which may lead to a more randomized sampling within the embedding space. Conversely, with the prompt in place, the model can directly target specific action embeddings.
@@
Figure 7(b) shows that the absence of the prompt is more detrimental in environments such as the Bernoulli Bandit, where each decision has a direct impact on the final outcome due to the environment’s single-episode structure. However, in multi-episode environments such as the Darkroom, the early lack of action space information becomes less impactful over time, as the model eventually encounters all actions through the context. This divergence highlights the prompt’s critical role in enabling the model to make informed choices in environments where each option has immediate and lasting consequences. Note that, as one can see in Figure 7(c), the impact of action set prompt ablation on the performance gets more pronounced as the number of arms grows.
@@
Contrastive Loss: In principle, the prediction of action embedding can be facilitated by various ways. Here, we consider the performance of a model that is asked to directly copy the corresponding embedding from its context to the output. To achieve it, we use mean squared error (MSE) loss instead of contrastive loss. In this case, probabilistic interpretation becomes less meaningful, which is why the nearest neighbor of the predicted vector is chosen as an action index, without sampling. The tuned hyperparameters can be found in Table 5.
Figure 9: Evaluating the Effectiveness of Contrastive vs. MSE Loss: This figure presents a comparative analysis of contrastive loss and mean squared error (MSE) loss on the training efficiency and performance metrics of our model. All results are averaged across various action sets within their respective environments and are based on data from five distinct training seeds. Performance on the y-axis represents the normalized regret for the Bernoulli Bandit environment, and the success rate for Darkroom. (a) depicts the impact on the model’s ability to explore: in the Bernoulli Bandit environment, the model trained with MSE loss exhibits a limited exploration range, focusing on only a small subset of the available actions. Conversely, on the Darkroom environment there is a slight increase in the number of actions attempted. Regardless of how the MSE loss affects the amount of actions used, the final performance of the model on both environments drops significantly.
@@
Figure 8(a) illustrates a significant decline in the variety of actions attempted by the model under this configuration within the bandit environment, showing that the model concentrates only on a fraction of the action set. Though the final regrets for this model are far from random, as depicted in Figure 8(b), we point out that this result is because the model skips exploration phase and goes directly to an exploitation of suboptimal actions. That allows it not to waste time on even more suboptimal ones via exploration. Appendix D shows the in-context curves of both models, providing more insight into MSE’s effect on model quality.
@@
Conversely, on the Darkroom environment, the loss substitution leads to a marginal increase in the amount of attempted actions. However, as one can see in Figure 8(b), this does not lead to increased performance. In fact, the performance of this model on the Darkroom environment is similar to random behavior.
@@
In summary, although the employed neural network is capable of learning an improvement operator and demonstrating ICL capabilities, the implemented loss function plays a crucial role in the success of this learning process. We emphasize the importance of our design choice to use contrastive loss by showing that a more naive approach of directly copying the action embeddings, as incentivized by MSE loss, results in underperformance.
@@ 7 Conclusion
In our work, we have introduced a new architecture that extends Algorithm Distillation (AD) for environments with variable action spaces, achieving invariance to their structure and size. Our approach consists of discarding the last linear layer, granting invariance to the action space structure, and making the model infer action semantics from the context, preparing it for the introduction of novel actions. We have demonstrated Headless-AD’s capability to generalize across new action spaces on a set of environments. We also observed its performance gains over vanilla AD, especially on larger action spaces. We hypothesize that this is due to the augmentation of the dataset by random embeddings, which was shown to improve the generalization abilities of agents (Kirsch et al., 2023). Headless-AD marks the progress toward versatile foundational models in RL, ones capable of operating across an expanded range of environments. We hope that Headless-AD inspires further development of models that can adapt to any action space beyond discrete ones.
8 Broader Impact
The work presented in this paper is aimed at advancing the field of Machine Learning. There can be many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ B.1 Transformers in Reinforcement Learning
The incorporation of transformers as a sequence modeling tool in RL began with the Decision Transformer (DT) by (Hu et al., 2022), which trains autoregressively on offline datasets of state, action, and return-to-go tuples. Unlike conventional RL, which focuses on return maximization, DT generates appropriate actions during inference by conditioning on specified return-to-go values. The Trajectory Transformer by Janner et al. (2021) introduces an alternative approach using beam search to bias trajectory samples based on future cumulative rewards. Building on this, the Multi-Game Decision Transformer (MGDT) by Lee et al. (2022) improves upon DT with enhanced transfer learning capabilities for new games, eliminating the need for manual return specification. Similarly, the Switch Trajectory Transformer by (Lin et al., 2022) expands on the Trajectory Transformer to facilitate multi-task training. However, when transitioning to novel tasks, these approaches require fine-tuning the model.
@@ B.3 In-Context Learning in RL
In-Context Learning (ICL) is an ability of a pretrained model to adapt and perform a new task given a context with examples $\{x_{i},f(x_{i})\}^{n}$, where $f$ is a function that gives ground truth targets (Brown et al., 2020; Wang et al., 2023). Previous work on ICL (Mirchandani et al., 2023) shows that Large Language Models operate as General Pattern Machines and are able to complete, transform and improve token sequences, even when the sequences consist of randomly sampled tokens. This ability supports our design choice to randomly encode the actions of agents. The research on Transformer Circuits (Elhage et al., 2021) gives evidence of Transformers’ ability to copy tokens, either literally or on a more abstract level, explaining their ICL capabilities. This is particularly useful for our model, which explicitly incentivizes an abstract copying of the action embeddings.
@@
Efforts to blend In-Context Learning (ICL) abilities of transformers with reinforcement learning (RL) are gaining traction, promising to create adaptable RL agents for real-world scenarios with varying conditions. Attempts to transfer ICL features to RL include Prompt-Based Decision Transformers (Xu et al., 2022), which leverage task-specific demonstration datasets as prompts, exhibiting strong few-shot learning without weight updates. Laskin et al. (2022) introduced Algorithm Distillation (AD) that trains a policy improvement operator using data from agent-environment interactions of learning RL algorithm, predicting the next action autoregressively. A key strategy here is across-episode training for capturing policy improvement. Lee et al. (2023) developed Decision Pretrained Transformer (DPT) using supervised training with unordered environmental interactions as context to predict optimal actions. Under certain conditions, DPT can be proved to implement posterior sampling, resulting in near-optimal exploration. However, DPT’s reliance on an optimal policy during training, not always available, is a limitation, similar to AD’s constraint on action space structure, restricting their use as fully generalist agents. Lin et al. (2023) provided theoretical analysis of AD and DPT, offering guarantees about learning from base algorithms. Zisman et al. (2023) propose a method for distilling the improvement operator from a demonstrator whose actions are initially noisy, with the noise level decreasing progressively throughout the data collection phase. This approach simplifies the data generation compared to AD as it does not require logging the training process of RL models and relaxes DPT’s limitation because the demonstrator may be suboptimal. Kirsch et al. (2023) demonstrated that data augmentation, through random projection of observations and actions, enhances task distribution and generalization on new domains, boosting ICL capabilities. Raparthy et al. (2023) study the effect of model size and dataset properties on the success of in-context learning. Other research, acknowledging Transformers’ limitations when it comes to long sequences, explores alternative models such as S4 (Gu et al., 2021). Lu et al. (2023) adapted S4 for RL, showing that it is capable of surpassing LSTM in performance and Transformers in runtime, indicating potential in RL ICL applications.
@@ B.4 Discarding the Linear Layer
Wolpertinger by Dulac-Arnold et al. (2015) also employs a similar approach of removing the linear head to improve the train and inference speeds. They suggest associating each action with an embedding and performing the training in a continuous action space. A specific action index is chosen as a nearest neighbor of the predicted embedding, and its effect is treated as environment dynamics. Most importantly, this nearest neighbor selection process can be optimized using an approximate algorithm. This leads to logarithmic time complexity, which is particularly advantageous in environments with large action spaces. Thus, we believe that Headless-AD is also capable of performing well on large action spaces, benefiting from runtime gains of approximate nearest neighbor lookup. However, Wolpertinger’s usage of fixed action embeddings prevents introduction of new actions, highlighting the significance of Headless-AD’s usage of randomized action embeddings.
@@ Appendix C Algorithm Distillation on Permuted Train Sets
Figure 10: AD Performance on Permuted Actions. This figure shows the models’ success rates on Darkroom environment, averaged over $5$ seeds. We evaluated AD’s ability to adapt to action sets with varied semantics by training it on a permuted dataset. Except for this, the training and testing are the same as in the vanilla setting and can be found in Section 4.3. Contrary to expectations, this tailored training did not enhance performance when compared to the AD model trained on standard datasets.
Appendix D MSE-Headless-AD In-Context Curves on Benroulli Bandit
Figure 11: MSE-Headless-AD In-Context Curves on Benroulli Bandit: Though Figure 8(b) shows that a variant of Headless-AD with MSE loss has a far from random regret, this graph shows that this result is due to the exploitation of several actions and the lack of exploration. We make this conclusion by observing that the shape of MSE-Headless-AD’s training curve is a straight line, showing the evidence that no learning is being present. All curves are averaged over $5$ seeds.
Appendix E Limitations and Future Work
Appendix F Sampling of Orthonormal Vectors
@@
Appendix G Algorithms’ Training Times
Table 1: Algorithm Training Times. Here we list the training times (in hours) of Headless-AD and AD on each environment. All experiments were performed on A100 GPUs. Note that times here include both training and evaluation steps. Though Headless-AD requires more time for completion, we point out that it is evaluated on more environment sets compared to AD.
@@ Appendix G Algorithms’ Training Times
Appendix H Model Hyperparameters
@@
Appendix I Code Sample
from typing import Tuple
[…]
import torch.nn as nn
[…]
def __init__(
self,
action_embed_dim: int,
num_actions: int,
%**** icml2024.tex Line 950 ****device: str,
super().__init__()
self.action_embed_dim = action_embed_dim
self.num_actions = num_actions
self.device = device
[…]
self.action_map.requires_grad = False
[…]
def regenerate(self) -> None:
torch.nn.init.orthogonal_(self.action_map, gain=1)
[…]
def __call__(self, actions: torch.Tensor) -> torch.Tensor:
embeds = self.action_map[actions]
[…]
def get_action(
self,
embeds: torch.Tensor,
) -> Tuple[torch.Tensor, torch.Tensor]:
%**** icml2024.tex Line 975 ****sims = embeds @ self.action_map.T
[…]
actions_sample, actions_argmax = dist.sample(), dist.probs.argmax(-1)
[…]
def _get_action_map_as_context(self, batch_size: int):
actions = torch.tile(self.action_map.unsqueeze(0), (batch_size, 1, 1))
[…]
act_mapper.regenerate()
action_embeds = act_mapper(actions)
[…]
states=states,
act_emb=action_embeds,
%**** icml2024.tex Line 1000 ****rewards=rewards,
actions_list=actions_list,
[…]
Figure 12: Code that demonstrates the Headless-AD training procedure. Note that this snippet is intended for illustration purposes only. The complete code can be found in Headless-AD’s repository, which will be open-sourced after the review.

### v5

@@ 5 Ablations
The hyperparameters tuned for each ablation can be found in Appendix J.
5.1 Action Set Prompt
@@
We assessed the impact of omitting the action embedding enumeration from the model context.
@@
Table 1 reveals that removing the action set prompt did not significantly alter the number of unique actions attempted by each model. This suggests that models without the prompt continue to sample a diverse range of actions, comparable to the original setup. However, the presence of the prompt did affect the performance. We hypothesize that, without the action set prompt, the model lacked explicit knowledge of the action space, which may have led to a more randomized sampling within the embedding space. Conversely, with the prompt in place, the model could directly target specific action embeddings.
@@
The absence of the prompt was more detrimental in the Bernoulli Bandit environment, where each decision has a direct impact on the final outcome due to the environment’s single-episode structure. However, in a multi-episode environment such as the Darkroom, the early lack of action space information becomes less impactful over time, as the model eventually encounters all actions through the context. This divergence highlights the prompt’s critical role in enabling the model to make informed choices in environments where each option has immediate and lasting consequences. Note that the impact of action set prompt ablation on the performance got more pronounced as the number of arms grows.
5.2 Contrastive Loss
@@
In principle, the prediction of action embedding can be facilitated by various ways. Here, we considered the performance of a model that was asked to directly copy the corresponding embedding from its context to the output. To achieve it, we used mean squared error (MSE) loss instead of contrastive loss.
In this case, probabilistic interpretation became less meaningful, which is why the nearest neighbor of the predicted vector was chosen as an action index, without sampling.
@@
Table 1 illustrates a significant decline in the variety of actions attempted by the model under this configuration within the bandit environment, showing that the model concentrated only on a fraction of the action set. Though the final regrets for this model were far from random, we point out that this result is because the model skipped exploration phase and went directly to an exploitation of suboptimal actions. That allowed it not to waste time on even more suboptimal ones via exploration. Appendix G shows the in-context curves of both models, providing more insight into MSE’s effect on model quality.
@@
Conversely, on the Darkroom environment, the loss substitution led to a marginal increase in the amount of attempted actions. However, this did not lead to an increased performance. In fact, the performance of this model on the Darkroom environment was similar to a random behavior.
@@
In summary, although the employed neural network architecture was capable of learning an improvement operator and demonstrating ICL capabilities, the implemented loss function played a crucial role in the success of this learning process. We emphasize the importance of our design choice to use contrastive loss by showing that a more naive approach of directly copying the action embeddings, as incentivized by MSE loss, resulted in underperformance.
5.3 Orthonormal Action Embeddings
In this ablation study, we underscored the significance of employing orthonormal vectors for action embeddings by contrasting them with vectors derived from a standard normal distribution. Our preference for the former stemmed from their ability to simplify the approximation of action probabilities because of their property of independence. Unlike embeddings from a standard normal distribution, orthonormal vectors ensure that assigning a probability weight to one vector does not unintentionally influence another (we illustrate it in Appendix K). This concept echoes the principle of superposition, observed when models incorporate more features than available dimensions, leading to feature interference (Elhage et al., 2022).
Table 1 illustrates the consequences of using linearly dependent action embeddings, manifesting as diminished performance across Bernoulli bandits and Darkroom scenarios. Notably, these results bear a striking resemblance to ”Prompt Ablation”, with both indicating a slight performance dip in Darkroom, a more noticeable decline in Bernoulli Bandits, and a general downtrend as the number of bandit arms grows. This parallel underscores a shared objective between these design choices: refining the model’s capacity for precise action selection.
@@ 7 Conclusion
In our work, we have introduced a new architecture that extends Algorithm Distillation (AD) for environments with variable action spaces, achieving invariance to their structure and size. Our approach consists of discarding the last linear layer, granting invariance to the action space structure, and making the model infer action semantics from the context, preparing it for the introduction of novel actions. We demonstrated Headless-AD’s capability to generalize across new action spaces on a set of environments. We also observed its performance gains over vanilla AD, especially on larger action spaces. We hypothesize that this is due to the augmentation of the dataset by random embeddings, which was shown to improve the generalization abilities of agents (Kirsch et al., 2023). Headless-AD marks the progress toward versatile foundational models in RL, ones capable of operating across an expanded range of environments. We hope that Headless-AD inspires further development of models that can adapt to any action space beyond discrete ones.
Impact Statement
Strong safeguards are necessary to prevent unauthorized users from manipulating the model, such as adding harmful action embeddings that could lead to negative outcomes.
Another concern is how the model handles out-of-distribution data. If the model encounters a new action that is significantly different from the training actions, it may take a while to understand its effects. Since our model learns by trying out actions, there is a risk it might perform harmful actions before learning they are inappropriate.
Any application of Headless-AD in real-life scenarios should be aware of these potential risks.
@@ B.1 Transformers in Reinforcement Learning
The incorporation of transformers as a sequence modeling tool in RL began with the Decision Transformer (DT) by (Hu et al., 2022), which is trained autoregressively on offline datasets of state, action, and return-to-go tuples. Unlike conventional RL, which focuses on return maximization, DT generates appropriate actions during inference by conditioning on specified return-to-go values. The Trajectory Transformer by Janner et al. (2021) is an alternative approach using beam search to bias trajectory samples based on future cumulative rewards. Building on this, the Multi-Game Decision Transformer (MGDT) by Lee et al. (2022) improves upon DT with enhanced transfer learning capabilities for new games, eliminating the need for manual return specification. Similarly, the Switch Trajectory Transformer by (Lin et al., 2022) expands on the Trajectory Transformer to facilitate multi-task training. However, when transitioning to novel tasks, these approaches require fine-tuning the model.
@@ B.3 In-Context Learning in RL
In-Context Learning (ICL) is an ability of a pretrained model to adapt and perform a new task given a context with examples $\{x_{i},f(x_{i})\}^{n}$, where $f$ is a function that gives ground truth targets (Brown et al., 2020; Wang et al., 2023). Previous work on ICL (Mirchandani et al., 2023) showed that Large Language Models operate as General Pattern Machines and are able to complete, transform and improve token sequences, even when the sequences consist of randomly sampled tokens. This ability supports our design choice to randomly encode the actions of agents. The research on Transformer Circuits (Elhage et al., 2021) gives evidence of Transformers’ ability to copy tokens, either literally or on a more abstract level, explaining their ICL capabilities. This is particularly useful for our model, which explicitly incentivizes an abstract copying of the action embeddings.
@@
Efforts to blend In-Context Learning (ICL) abilities of transformers with reinforcement learning (RL) are gaining traction, promising to create adaptable RL agents for real-world scenarios with varying conditions. Attempts to transfer ICL features to RL include Prompt-Based Decision Transformers (Xu et al., 2022), which leverage task-specific demonstration datasets as prompts, exhibiting strong few-shot learning without weight updates. Laskin et al. (2022) introduced Algorithm Distillation (AD) that trains a policy improvement operator using data from agent-environment interactions of learning RL algorithm, predicting the next action autoregressively. A key strategy here is across-episode training for capturing policy improvement. Lee et al. (2023) developed Decision Pretrained Transformer (DPT) using supervised training with unordered environmental interactions as context to predict optimal actions. Under certain conditions, DPT can be proved to implement posterior sampling, resulting in near-optimal exploration. However, DPT’s reliance on an optimal policy during training, not always available, is a limitation, similar to AD’s constraint on action space structure, restricting their use as fully generalist agents. Lin et al. (2023) provided theoretical analysis of AD and DPT, offering guarantees about learning from base algorithms. Wang et al. (2024) also provided a theoretical analysis showing that transformers may implement several RL algorithm in-context. Zisman et al. (2023) propose a method for distilling the improvement operator from a demonstrator whose actions are initially noisy, with the noise level decreasing progressively throughout the data collection phase. This approach simplifies the data generation compared to AD as it does not require logging the training process of RL models and relaxes DPT’s limitation because the demonstrator may be suboptimal. Kirsch et al. (2023) demonstrated that data augmentation, through random projection of observations and actions, enhances task distribution and generalization on new domains, boosting ICL capabilities. Raparthy et al. (2023) study the effect of model size and dataset properties on the success of in-context learning. Other research, acknowledging Transformers’ limitations when it comes to long sequences, explores alternative models such as S4 (Gu et al., 2021). Lu et al. (2023) adapted S4 for RL, showing that it is capable of surpassing LSTM in performance and Transformers in runtime, indicating potential in RL ICL applications.
@@ B.4 Discarding the Linear Layer
Wolpertinger by Dulac-Arnold et al. (2015) also employs a similar approach of removing the linear head to improve the train and inference speeds. The authors suggest associating each action with an embedding and performing the training in a continuous action space. A specific action index is chosen as a nearest neighbor of the predicted embedding, and its effect is treated as environment dynamics. Most importantly, this nearest neighbor selection process can be optimized using an approximate algorithm. This leads to logarithmic time complexity, which is particularly advantageous in environments with large action spaces. Thus, we believe that Headless-AD is also capable of performing well on large action spaces, benefiting from runtime gains of approximate nearest neighbor lookup. However, Wolpertinger’s usage of fixed action embeddings prevents introduction of new actions, highlighting the significance of Headless-AD’s usage of randomized action embeddings.
@@ Appendix C Algorithm Distillation on Permuted Train Sets
Figure 8: AD Performance on Permuted Actions: This figure shows the models’ success rates on Darkroom environment, averaged over $5$ seeds. We evaluated AD’s ability to adapt to action sets with varied semantics by training it on a permuted dataset. Except for this, the training and testing are the same as in the vanilla setting and can be found in Section 4.3. Contrary to expectations, this tailored training did not enhance performance when compared to the AD model trained on standard datasets.
Appendix D Across-Environment Generalization
We evaluated Headless-AD’s ability to generalize skills learned in one domain to another. We trained Headless-AD on Contextual Bandits and subsequently assessed its performance on Bernoulli Bandits. To align with the Contextual Bandit setting, we introduced a random vector to serve as the state in the Bernoulli Bandit environment, effectively transforming it into a single-context Contextual Bandit scenario with Bernoulli-distributed rewards.
Figure 9 presents the results of these cross-domain experiments. Headless-AD not only adapted to the new domain but also maintained a satisfactory performance level, highlighting the model’s capacity for cross-environment application.
Figure 9: Evaluation of Across-Environment Generalization: This graph illustrates the performance of the Headless-AD model, which was initially trained on a Contextual Bandit setting and then evaluated on a Bernoulli Bandit environment. The experiment aimed to assess the model’s generalization capabilities across novel environments. As shown, Headless-AD achieves a decent performance across various configurations with different numbers of arms, demonstrating its potential for across-environment usage.
Appendix E Visual Darkroom
We adapted the Darkroom environment to produce high-dimensional visual observations to demonstrate Headless-AD’s capability to generalize to new action spaces in more complex settings. In this modified version, the grid is visually represented, with the agent’s position indicated in red (see Figure 10). This change introduces a more complex observation space. For the model to handle these high-dimensional visual observations, we integrated a convolutional network that converts the visual input into an embedding of size ’token_dim’.
Figure 10: Observation in the Visual Darkroom Environment.
Both Headless-AD and AD were configured with the identical hyperparameters previously applied in the Darkroom experiments. The results at Figure 11 show that Headless-AD maintains its superior performance in this more complex observational setting, whereas AD experiences a notable drop in performance. This discrepancy likely stems from our decision to retain the original hyperparameters without tuning them for the new environment. Based on our experience, AD’s performance is particularly sensitive to hyperparameter settings.
Figure 11: Visual Darkroom: Darkroom environment was modified to emit visual observations. Headless-AD consistently outperforms AD in this more complex observational setting.
Appendix F Darkroom. Alternative Split
A random split of actions in Darkroom environment may lead to a data leakage as multiple action sequences result in the same endpoint. To address this, we conducted an experiment with distinct, non-overlapping action sets for training and testing. We trained Headless-AD with actions that cause movements of 00, $1$, and $3$ cells, and tested on movements of $2$ cells, introducing scenarios not encountered during training.
Figure 12 shows that although Headless-AD experienced a slight decrease in performance on the test action set, it still remained significantly above the random level and outreaching the AD-level, underscoring its capability to generalize beyond the training action space. This experiment intends to make the capabilities and limitations of Headless-AD clearer.
Figure 12: Darkroom. Alternative Split: In this experiment the actions in Darkroom were split into non-overlapping in terms of the distance of the endpoint from the initial position. Train split consisted of actions with length 00, $1$ and $3$, test split - $2$. Headless-AD, though with a slight drop, maintained its performance on levels seen during the previous experiment with a random split.
loh
loh
Appendix G MSE-Headless-AD In-Context Curves on Bernoulli Bandit
Figure 13: MSE-Headless-AD In-Context Curves on Benroulli Bandit: Though Table 1 shows that a variant of Headless-AD with MSE loss has a far from random regret, this graph shows that this result is due to the exploitation of several actions and the lack of exploration. We make this conclusion by observing that the shape of MSE-Headless-AD’s training curve is a straight line, showing the evidence that no learning is being present. All curves are averaged over $5$ seeds.
Appendix H Sampling of Orthonormal Vectors
@@
Appendix I Algorithms’ Training Times
Table 2: Algorithm Training Times: Here we list the training times (in hours) of Headless-AD and AD on each environment. All experiments were performed on A100 GPUs. Note that times here include both training and evaluation steps. Though Headless-AD requires more time for completion, we point out that it is evaluated on more environment sets compared to AD.
@@ Appendix I Algorithms’ Training Times
Appendix J Model Hyperparameters
@@
Table 7: Headless-AD’s Environment-Specific Hyperparameters for Action Embeddings Ablation: For certain instances, hyperparameters underwent optimization within the specified ranges in the Sweep Values column, utilizing the Bayesian search method facilitated by the wandb sweep tool (Biewald, 2020). This process was employed to identify the optimal set of hyperparameters for enhanced performance and fair comparisons.
Hyperparameter Bernoulli Bandit Darkroom Sweep Values
Number of Layers 4 4 [1,2,4,6,8]
Number of Heads 64 64 [2,4,8,16,32,64
Model Dim. 512 512 [128, 256, 512, 1024, 2048]
Sequence Length 300 100
Learning Rate 1.1e-03 6.6e-04 [1e-5, 1e-2]
Weight Decay 1.2e-03 9.1e-04 [1e-6, 5]
$\beta_{1}$ 0.94 0.6 [0, 1]
Attention Dropout Rate 0.62 0.53 [0, 1]
Dropout Rate 0.17 3.0e-02 [0, 1]
In-Context Episodes 300 1000
Action Selection sample sample
Appendix K Linear Dependence of Different Types of Action Embeddings
We illustrate how the probability mass can unintentionally be put on unwanted actions through an analysis of cosine similarities between action embeddings, comparing orthonormal vectors to those from a standard normal distribution. Figure 14 demonstrates non-zero cosine similarities for the standard normal distribution, indicating that a vector perfectly aligned with one action embedding may erroneously attribute non-zero probabilities to other actions.
Figure 14: These plots illustrate the pairwise cosine similarities between action embeddings for orthonormal versus standard normal distributions. (a) In a $128$-dimensional space with $10$ actions, orthonormal embeddings exhibit perfect decorrelation, whereas standard normal embeddings display non-zero similarities among them. (b) With $10$ actions in a $16$-dimensional space, the similarity between embeddings increases as the dimensionality of the space decreases, indicating denser correlations. (c) For $10$ actions in an $8$-dimensional space, despite the action count exceeding the space’s dimensionality, orthonormal vectors (referenced in Appendix: Orthonormal Vectors) maintain lower similarities compared to those from a standard normal distribution, emphasizing the robustness of orthonormal embeddings in constrained dimensions.
Appendix L Code Sample
{listing}
[h!]
{minted} [ fontsize=, fontfamily=tt ]python from typing import Tuple import torch import torch.nn as nn class ActionMapper(nn.Module): def __init__( self, action_embed_dim: int, num_actions: int, device: str, ): super().__init__() self.action_embed_dim = action_embed_dim self.num_actions = num_actions self.device = device self.action_map = torch.empty((num_actions, action_embed_dim), device=device) self.action_map.requires_grad = False @torch.no_grad() def regenerate(self) -¿ None: torch.nn.init.orthogonal_(self.action_map, gain=1) @torch.no_grad() def __call__(self, actions: torch.Tensor) -¿ torch.Tensor: embeds = self.action_map[actions] return embeds @torch.no_grad() def get_action( self, embeds: torch.Tensor, ) -¿ Tuple[torch.Tensor, torch.Tensor]: sims = embeds @ self.action_map.T dist = torch.distributions.Categorical(logits=sims) actions_sample, actions_argmax = dist.sample(), dist.probs.argmax(-1) return actions_sample, actions_argmax @torch.no_grad() def _get_action_map_as_context(self, batch_size: int): actions = torch.tile(self.action_map.unsqueeze(0), (batch_size, 1, 1)) return actions act_mapper = ActionMapper(…) for states, actions, rewards in dataloader: act_mapper.regenerate() action_embeds = act_mapper(actions) actions_list = act_mapper._get_action_map_as_context(batch_size=states.shape[0]) pred = model( states=states, act_emb=action_embeds, rewards=rewards, actions_list=actions_list, ) act_sample, act_mode = act_mapper.get_action(pred)
Code that demonstrates the Headless-AD training procedure. Note that this snippet is intended for illustration purposes only. The complete code can be found in Headless-AD’s repository.